How do you solve-
A squared + 15 = 2A squared - 65?

Answer by dabanfield(803)

A squared + 15 = 2A squared - 65?
a^2 + 15 = (2a)^2 - 65
a^2 + 15 = 4a^2 - 65
3a^2 = 80
a^2 = 80/3
a = sqrt(80)
a = sqrt(16*5)
a = sqrt(16)*sqrt(5)
a = 4*sqrt(5)